### v2

# Estimating Rate-Distortion Functions Using the Energy-Based Model

Abstract
The rate-distortion (RD) theory is one of the key concepts in information theory, providing theoretical limits for compression performance and guiding the source coding design, with both theoretical and practical significance. The Blahut-Arimoto (BA) algorithm, as a classical algorithm to compute RD functions, encounters computational challenges when applied to high-dimensional scenarios. In recent years, many neural methods have attempted to compute high-dimensional RD problems from the perspective of implicit generative models. Nevertheless, these approaches often neglect the reconstruction of the optimal conditional distribution or rely on unreasonable prior assumptions. In face of these issues, we propose an innovative energy-based modeling framework that leverages the connection between the RD dual form and the free energy in statistical physics, achieving effective reconstruction of the optimal conditional distribution. The proposed algorithm requires training only a single neural network and circumvents the challenge of computing the normalization factor in energy-based models using the Markov chain Monte Carlo (MCMC) sampling. Experimental results demonstrate the significant effectiveness of the proposed algorithm in estimating high-dimensional RD functions and reconstructing the optimal conditional distribution.
Index Terms:
Rate-distortion function, energy-based models, neural networks
I Introduction
The rate-distortion (RD) theory, first introduced by Shannon in [1, 2], is one of the fundamental concepts in information theory. It characterizes the minimum amount of information (rate) required to represent a source while maintaining a specified level of distortion. By establishing theoretical limits on compression performance, the RD theory not only guides the development of practical source coding schemes [3, 4] but also enables the evaluation of their capabilities. Therefore, solving the RD problem is both a theoretically significant and practically essential endeavor.
It is known that the classical Blahut-Arimoto (BA) algorithm [5] can efficiently solve the RD problem with discrete source distributions. However, when applied to continuous cases, the algorithm necessitates discretization, which results in an exponential increase in computational complexity[6, 7]. Therefore, the BA algorithm is not always applicable for continuous and high-dimensional scenarios.
In recent years, several novel approaches have been proposed to solve the high-dimensional RD problem. For instance, the authors of [8] provide sandwich bounds for the RD function, where the upper bound is obtained utilizing the $\beta$-VAE [9] architecture and has been validated to be tight for high-dimensional scenarios. Meanwhile, [7] represents the reproduction distribution by a mapping and optimizes the dual function via neural parameterization. Another work [10] introduces a neural-network-free algorithm, using Wasserstein gradient descent in optimal transport to optimize the distribution.
However, while the dual function framework utilized in [7, 10] allows for a more compact formulation, a limitation of these works is that they consider only the optimization of the marginal distribution, neglecting the reconstruction of the conditional distribution. This is noteworthy since the conditional distribution is not only inherent to the formulation of the RD problem but also provides valuable insights for compression. Although the method in [8] can obtain both the marginal distribution and the conditional distribution, it relies on an assumption about the prior of the reproduction data, which is not always reasonable. As a result, effective and accurate reconstruction of the conditional distribution still remains unaddressed.
Existing neural methods [7, 8] estimating the RD function primarily model probability distribution through the lens of implicit generative models. However, with the recent breakthroughs in diffusion models [11] and score-based models [12], the energy-based modeling framework has emerged as a compelling alternative to probabilistic modeling. Notably, a recent work [13] approaches the RD problem from an energy perspective, finding that the optimal conditional distribution and the optimal marginal distribution are linked via the Boltzmann distribution. The authors of [13] innovatively employ a single network to represent both distributions. Nevertheless, this method does not fully leverage the properties of the RD function. It approximates the mutual information in the original RD Lagrangian with an additional neural network [14], increasing the estimation error and computational complexity.
In this paper, we propose a more concise algorithm for estimating RD functions using the energy-based model. Different from [13], we recognize that the dual form of the RD function corresponds to the free energy in statistical physics [15], based on which our approach follows. Specifically, by substituting the marginal distribution, as represented by the energy-based model, into the dual function, we obtain the optimization objective. To avoid the challenge posed by the intractable normalization factor in energy-based models, we compute the gradient of the objective and estimate it using Langevin Markov chain Monte Carlo (Langevin MCMC) methods [16, 17], which enable practical training of the energy function through gradient-based optimization. Furthermore, we analyze the asymptotic property of the proposed estimation algorithm and elucidate its connection with classical energy-based models. Finally, numerical experiments show the effectiveness of the proposed algorithm.
II Preliminaries
II-A Energy-Based Models
The energy-based model (EBM) [18] is a framework for modeling the probability density of variables with an energy function. Specifically, let $x$ denote a random variable in the continuous space $\mathcal{X}$ and $E_{\theta}(x):\mathcal{X}\rightarrow\mathcal{R}$ represent an energy function parameterized by a neural network. The probability density of $x$ is given by the Boltzmann distribution:
$p_{\theta}(x)=\frac{e^{-E_{\theta}(x)}}{Z_{\theta}},\vskip-7.0pt$
where $Z_{\theta}=\int_{\mathcal{X}}e^{-E_{\theta}(x)}\mathrm{d}x$ is the normalization factor.
The key advantage of energy-based models lies in their ability to efficiently compute the gradient of the log-probability (also known as the score) with respect to $x$:
$\nabla_{x}\log p_{\theta}(x)=-\nabla_{x}E_{\theta}(x).\vskip-5.0pt$
Then a sample of $x$ can be obtained through the Langevin MCMC [16, 17]. The update rule is given by
$x_{k}=x_{k-1}-\frac{\epsilon^{2}}{2}\nabla_{x}E_{\theta}(x_{k-1})+\epsilon z_{k},\vskip-2.0pt$
where $\epsilon$ is the step size, and $z_{k}\sim\mathcal{N}(0,1)$ is a noise term. As $\epsilon\rightarrow 0$ and the number of steps $K\rightarrow+\infty$, the distribution of $x_{K}$ is guaranteed to converge to $p_{\theta}(x)$.
The energy function $E_{\theta}(x)$ can be trained using maximum likelihood estimation (MLE)[19, 20]. Specifically, let $p_{\text{data}}(x)$ be the real distribution of the data, and then the target is to minimize the expectation of negative log-likelihood function
$\mathbb{E}_{x\sim p_{\text{data}}(x)}\left[-\log p_{\theta}(x)\right].\vskip-2.0pt$
The function itself is hard to estimate as $Z_{\theta}$ is intractable. Nevertheless, its gradient can be simplified by
$\nabla_{\theta}\mathbb{E}_{x\sim p_{\text{data}}(x)}\left[-\log p_{\theta}(x)\right]$
$=$ $-\mathbb{E}_{x\sim p_{\text{data}}(x)}\left[-\nabla_{\theta}E_{\theta}(x)\right]+\nabla_{\theta}\log Z_{\theta}$
$=$ $\mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\nabla_{\theta}E_{\theta}(x)\right]-\mathbb{E}_{x\sim p_{\theta}(x)}\left[\nabla_{\theta}E_{\theta}(x)\right].$
Thus, it can be estimated using Monte Carlo methods to perform gradient descent.
II-B Rate-Distortion Theory
The RD function characterizes the trade-off between the compression rate and reconstruction fidelity in lossy source coding. Given a memoryless source $X\sim P_{X}$ supported on the alphbet $\mathcal{X}$ with reproduction $Y\sim\mathcal{Y}$, the RD function is defined as
$R(D):=\inf_{P_{Y\mid X}:\mathbb{E}_{P_{X,Y}}[\rho(X,Y)]\leq D}I(X;Y),$ (1)
where $\rho:\mathcal{X}\times\mathcal{Y}\rightarrow[0,+\infty)$ is the distortion measure.
A classical numerical approach to computing the RD function is the Blahut-Arimoto (BA) algorithm [5]. This method addresses the RD problem in discrete scenarios by iteratively optimizing the following two variables:
[…]
where $\beta$ is a fixed Lagrangian multiplier. The algorithm works well for low-dimensional cases but becomes ineffective or even infeasible as the dimension increases[7].
By performing variational calculation on the Lagrangian of the RD function[21], the problem can be transformed into minimizing
$F(q)=-\int_{\mathcal{X}}\mathrm{d}xp(x)\log\int_{\mathcal{Y}}\mathrm{d}yq(y)e^{-\beta\rho(x,y)}.$ (2)
This formulation is flexible and has led to the development of neural network algorithms specifically designed for high-dimensional cases[7].
II-C Connection between Rate-Distortion Problem and Energy-Based Models
A key connection between energy-based models and the rate-distortion problem lies in the observation that the optimality condition of the RD problem, i.e.,
$p(y|x)=\frac{q(y)e^{-\beta\rho(x,y)}}{\int_{\mathcal{Y}}q(y)e^{-\beta\rho(x,y)}\mathrm{d}y},$ (3)
also takes the form of a Boltzmann distribution.
The authors of [13] demonstrated that it is feasible to use a single neural network to represent both distributions $q(y)$ and $p(y|x)$, as detailed in the following lemma.
Lemma 1.
In the RD problem, if the optimal marginal distribution $q(y)$ is represented by the energy function $E_{\theta}(y)$, the conditional distribution $p(y|x)$ can be represented by $E^{\prime}_{\theta}(x,y)=E_{\theta}(y)+\beta\rho(x,y)$, i.e.,
$q_{\theta}(y)=\frac{e^{-E_{\theta}(y)}}{\int_{\mathcal{Y}}e^{-E_{\theta}(y)}\mathrm{d}y},$
$p_{\theta}(y|x)=\frac{e^{-E^{\prime}_{\theta}(x,y)}}{\int_{\mathcal{Y}}e^{-E^{\prime}_{\theta}(x,y)}\mathrm{d}y}.$
However, the subsequent relaxation of mutual information in [13] increases the complexity of the algorithm. In this paper, we take Lemma 1 as a start point, and then investigate more intrinsic connections between the RD function and energy-based models, subsequently devising a more concise algorithm.
III Energy-Based Model of Rate-Distortion Function
In this section, we introduce the energy-based rate-distortion (EBRD) model. We present the derivation of the algorithm and discuss its relationship with the classical EBMs.
III-A Energy-Based Rate-Distortion Model
Noting that the dual function (2) of RD corresponds to the free energy in statistical physics [15], we adopt this duality as our optimization objective. By representing the distribution $q(y)$ in (2) using an energy-based model $E_{\theta}(y)$, the loss function to be minimized can be obtained as:
$\mathcal{L}(\theta)$ $=-\mathbb{E}_{P_{X}}\left[\log\mathbb{E}_{Q_{Y}^{\theta}}\left[e^{-\beta\rho(X,Y)}\right]\right]$
$=-\mathbb{E}_{P_{X}}\left[\log\int_{Y}e^{-E_{\theta}^{\prime}(X,y)}\mathrm{d}y-\log\int_{Y}e^{-E_{\theta}(y)}\mathrm{d}y\right].$
Similar to MLE, there is an intractable term $\int_{\mathcal{Y}}e^{-E_{\theta}(y)}\mathrm{d}y$ in $\mathcal{L}(\theta)$, which makes direct computation infeasible. Therefore, our goal is to compute its gradient with respect to $\theta$, as described in Proposition 4.
Proposition 1.
Let $P^{\theta}_{Y|X}$ be the conditional distribution represented by $E^{\prime}_{\theta}(x,y)=E_{\theta}(y)+\beta\rho(x,y)$, and $P^{\theta}_{Y}$ be the marginal distribution of the joint distribution $P_{X}P^{\theta}_{Y|X}$ over the space $\mathcal{Y}$. Then the gradient of the objective function $\mathcal{L}(\theta)$ with respect to the parameter $\theta$ is:
$\nabla_{\theta}\mathcal{L}(\theta)\!=\!\mathbb{E}_{P^{\theta}_{Y}}\!\left[\nabla_{\theta}E_{\theta}(Y)\right]\!-\!\mathbb{E}_{Q^{\theta}_{Y}}\!\left[\nabla_{\theta}E_{\theta}(Y)\right].$ (4)
Proof.
First, we expand the expression for the gradient of $\mathcal{L}(\theta)$ with respect to $\theta$:
$\nabla_{\theta}\mathcal{L}(\theta)$ $\!=\!-\nabla_{\theta}\mathbb{E}_{P_{X}}\!\!\left[\log\!\int_{\mathcal{Y}}e^{-E_{\theta}^{\prime}(X,y)}\mathrm{d}y\!-\!\log\!\int_{\mathcal{Y}}e^{-E_{\theta}(y)}\mathrm{d}y\right]$
$\!=\!-\mathbb{E}_{P_{X}}\!\!\left[\nabla_{\theta}\!\log\!\!\int_{\mathcal{Y}}\!e^{-E_{\theta}^{\prime}(X,y)}\mathrm{d}y\!-\!\nabla_{\theta}\!\log\!\!\int_{\mathcal{Y}}\!e^{-E_{\theta}(y)}\mathrm{d}y\right]\!.$
For the first term, applying the chain rule of differentiation, we obtain
$\nabla_{\theta}\log\int_{\mathcal{Y}}e^{-E_{\theta}^{\prime}(X,y)}\mathrm{d}y$ $=\frac{\int_{\mathcal{Y}}e^{-E_{\theta}^{\prime}(X,y)}(-\nabla_{\theta}E_{\theta}^{\prime}(X,y))\mathrm{d}y}{\int_{\mathcal{Y}}e^{-E_{\theta}^{\prime}(X,y)}\mathrm{d}y}$
$=\int_{\mathcal{Y}}p_{\theta}(y|x)(-\nabla_{\theta}E_{\theta}^{\prime}(X,y))\mathrm{d}y,$
where $\nabla_{\theta}E_{\theta}^{\prime}(X,y)=\nabla_{\theta}E_{\theta}(y)$ due to $E^{\prime}_{\theta}(x,y)=E_{\theta}(y)+\beta\rho(x,y)$. Thus,
$\nabla_{\theta}\log\int_{\mathcal{Y}}e^{-E_{\theta}^{\prime}(X,y)}\mathrm{d}y$ $=\int_{\mathcal{Y}}p_{\theta}(y|x)(-\nabla_{\theta}E_{\theta}(y))\mathrm{d}y$
$=-\mathbb{E}_{P^{\theta}_{Y|X}}\left[\nabla_{\theta}E_{\theta}(Y)\right].$
The second term can also be derived in a similar manner:
$\nabla_{\theta}\log\int_{\mathcal{Y}}e^{-E_{\theta}(y)}\mathrm{d}y$ $=\frac{\int_{\mathcal{Y}}e^{-E_{\theta}(y)}(-\nabla_{\theta}E_{\theta}(y))\mathrm{d}y}{\int_{\mathcal{Y}}e^{-E_{\theta}(y)}\mathrm{d}y}$
$=\int_{\mathcal{Y}}q_{\theta}(y)(-\nabla_{\theta}E_{\theta}(y))\mathrm{d}y$
$=-\mathbb{E}_{Q^{\theta}_{Y}}\left[\nabla_{\theta}E_{\theta}(Y)\right].$
By substituting the simplified two terms, the gradient can be further simplified as
$\nabla_{\theta}\mathcal{L}(\theta)$ $\!=\!\mathbb{E}_{P_{X}}\!\left[\mathbb{E}_{P^{\theta}_{Y|X}}\!\left[\nabla_{\theta}E_{\theta}(Y)\right]\!-\!\mathbb{E}_{Q^{\theta}_{Y}}\!\left[\nabla_{\theta}E_{\theta}(Y)\right]\right]$
$\!=\!\mathbb{E}_{P^{\theta}_{Y}}\!\left[\nabla_{\theta}E_{\theta}(Y)\right]\!-\!\mathbb{E}_{Q^{\theta}_{Y}}\!\left[\nabla_{\theta}E_{\theta}(Y)\right].$
Thus, the final formula (4) is obtained. ∎
In practice, one can approximate the expectation term in (4) using Langevin MCMC. The specific sampling steps are described as follows:
  1. 1.
    Sample $x_{1},x_{2},\dots,x_{N}$ from $P_{X}$, i.e., derive a random batch from the source dataset.
  2. 2.
    For each $x_{i}$, sample $y_{i}$ from the conditional distribution $P^{\theta}_{Y|X}$ via Langevin MCMC.
  3. 3.
    Sample $y^{\prime}_{1},y^{\prime}_{2},\dots,y^{\prime}_{N}$ from $Q^{\theta}_{Y}$ via Langevin MCMC.
Therefore, the formula (4) can be approximated as:
$\nabla_{\theta}\mathcal{L}(\theta)\!\approx\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}E_{\theta}(y_{i})-\frac{1}{N}\sum_{j=1}^{N}\nabla_{\theta}E_{\theta}(y^{\prime}_{j}).$
To this end, we have derived a computable approximate form for the gradient of the loss function $\mathcal{L}_{\theta}$. Therefore, we can use gradient descent to train the energy function $E_{\theta}(y)$. The detailed training procedure is summarized in Algorithm 1. Additionally, we describe in Appendix A how to estimate the $(R,D)$ pairs using the well-trained energy function $E_{\theta}$ and establish theoretical guarantees for the proposed estimation method.
Algorithm 1 Training Process of EBRD Model
0: Source distribution $P_{X}$; distortion measure $\rho(x,y)$; number of steps $K$, and step size $\epsilon$ for Langevin MCMC; sample batch size $N$; learning rate $\eta$ and hyper-parameter $\beta$.
0: Parameter $\theta$ of the energy function $E_{\theta}(y)$.
while $\theta$ does not converge do
  Sample $x_{1},x_{2},\ldots,x_{N}\sim P_{X}$
  Sample $y^{(0)}_{1},y^{(0)}_{2},\ldots,y^{(0)}_{N}\sim\mathcal{N}(0,I)$
  for $k=1,2,\ldots,K$ do
   Obtain $y^{(k)}_{i},i=1,2,\ldots,N$ through Langevin step:
$y^{(k)}_{i}=y^{(k-1)}_{i}-\frac{\epsilon^{2}}{2}\nabla_{y}E^{\prime}_{\theta}(x_{i},y^{(k-1)}_{i})+\epsilon z^{(k)}_{i},$
where $z^{(k)}_{i}\sim\mathcal{N}(0,I)$.
  Sample $y^{\prime(0)}_{1},y^{\prime(0)}_{2},\ldots,y^{\prime(0)}_{N}\sim\mathcal{N}(0,I)$
  for $k=1,2,\ldots,K$ do
   Obtain $y^{\prime(k)}_{j},j=1,2,\ldots,N$ through Langevin step:
$y^{\prime(k)}_{j}=y^{\prime(k-1)}_{j}-\frac{\epsilon^{2}}{2}\nabla_{y}E_{\theta}(y^{\prime(k-1)}_{j})+\epsilon z^{(k)}_{j},$
where $z^{(k)}_{i}\sim\mathcal{N}(0,I)$.
  Compute the approximate gradient:
$\nabla_{\theta}\mathcal{L}(\theta)\!\approx\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}E_{\theta}(y^{(K)}_{i})-\frac{1}{N}\sum_{j=1}^{N}\nabla_{\theta}E_{\theta}(y^{\prime(K)}_{j}).$
  Update $\theta$ by gradient descent $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}(\theta)$
Return:  $\theta$
III-B Relation with Classical Energy-Based Models
The proposed EBRD model and the classical EBMs are formally very similar. However, a notable difference lies in the approach used to compute the gradients of the network parameter $\theta$. The EBM computes the difference between the energy gradients of the original data and the generated samples. In contrast, the EBRD model computes the difference between the energy gradients of samples generated from the conditional distribution and those generated directly from the marginal distribution. Fig. 1 illustrates the sampling processes of the two models.
As $\beta\rightarrow+\infty$, the conditional distribution $p_{\theta}(y|x)$ converges to $\delta_{x}(y)$. In this case, it can be assumed that the samples generated from the conditional distribution are identical to the original samples, suggesting that the classical EBM and the EBRD model become asymptotically equivalent.
Figure 1: Comparison of sampling schemes between the classical EBM and the EBRD model.
IV Experimental Results and Discussions
This section evaluates the effectiveness of the proposed EBRD model. We first test the accuracy of estimating the RD function for scalar classical sources and vector Gaussian sources, and then visualize the reconstruction of conditional distributions on a two-dimensional Gaussian mixture model.
IV-A Cases with Scalar Classical Sources
This subsection estimates the RD function of two classical sources, i.e., the Gaussian source and the Laplacian source.
For the Gaussian source, we set $X\sim\mathcal{N}(0,1)$, and employ the squared Euclidean distance $\|x-y\|^{2}$ as the distortion measure. Its RD function is given by [22]:
$R_{\text{gaussian}}(D)=\left\{\begin{aligned} &\frac{1}{2}\log(\frac{1}{D}),&0\leq D\leq 1\\ &0,&D>1\end{aligned}\right..$
For the Laplacian source, the density function of $X$ is
$p(x)=\frac{1}{2}\exp\left(-|x|\right).$
[…]
$R_{\text{laplacian}}(D)=\left\{\begin{aligned} &\log(\frac{1}{D}),&0\leq D\leq 1\\ &0,&D>1\end{aligned}\right..$
[…]
Figure 2: Comparison between the theoretical result and the EBRD algorithm on Gaussian (left) and Laplacian (right) sources.
IV-B Evaluation for Vector Gaussian Sources
This subsection evaluates the performance of the proposed EBRD model on high-dimensional Gaussian sources.
Specifically, let $X$ follow the distribution $\mathcal{N}(0,\Sigma)$, where $\Sigma$ admits the decomposition $\Sigma=U\text{diag}(\sigma_{1}^{2},\sigma_{2}^{2},\ldots,\sigma_{d}^{2})U^{T}$ with $U$ being an orthogonal matrix. According to [22], its RD function is
$R(D)=\left\{\begin{aligned} &\sum_{i=1}^{d}\frac{1}{2}\log\frac{\sigma_{i}^{2}}{D_{i}},&0\leq D\leq\sum_{i=1}^{d}\sigma_{i}^{2}\\ &0,&D>\sum_{i=1}^{d}\sigma_{i}^{2}\end{aligned}\right..$
where $D_{i}\!=\!\min(\lambda,\sigma_{i}^{2})$, and $\lambda$ is chosen to satisfy the equality $\sum_{i=1}^{d}D_{i}=D$.
In the experiment, we randomly $U$ and set $\sigma_{i}=2^{-\frac{i}{10}}$. Fig. 3 presents the experimental results obtained by the EBRD algorithm for dimensions $d=2,5,10$ compared with the NERD algorithm[7]. It can be seen that both methods consistently obtain accurate results for $d=2$ and 5. As for $d=10$, the results from both algorithms are accurate at low to medium rates. However, for sufficiently high rates, errors emerge in both algorithms since the number of samples required for accurate computation grows exponentially with the rate, as explained in [7].
Figure 3: Comparison of EBRD and NERD algorithms with theoretical curves for vector Gaussian sources, with $d=2,5,10$ from left to right.
IV-C Visualization of Conditional Sampling
This subsection adopts a two-dimensional Gaussian mixture model to visualize the optimal conditional distribution.
This distribution is constructed as a mixture of three two-dimensional Gaussian distributions with equal weights, each having distinct means and identical variances. The density function of this distribution is
$p(x)=\sum_{i=1}^{3}\frac{1}{3}\mathcal{N}(\mu_{i},\Sigma_{i}).$
Specifically, the means of the three Gaussian components are placed equidistantly on a circle with a radius of 6. The variance of each component is set to 111. Samples are generated by first randomly selecting a component with equal probability and then drawing from the corresponding Gaussian distribution.
[…]
V Conclusion
In this work, we delve deeply into the connection between energy-based models and the RD problem, and propose an effective method to estimate the RD functions, i.e., the EBRD algorithm. The proposed algorithm is based on a reformulation of the RD problem by analogizing to the free energy in statistical physics, and then we derive a gradient descent algorithm similar to that in the classical EBMs. The proposed EBRD algorithm not only accurately estimates the RD curve but also successfully reconstructs the optimal conditional distribution. Moreover, numerical experiments validate the effectiveness of the EBRD algorithm.
As for the future work, we aim to improve the sampling scheme involved in the EBRD model to enhance the stability of computation. This refinement would allow the algorithm to handle larger-scale datasets, further expanding its applicability and potential impact.
Figure 4: 5000 sampled points from the Gaussian mixture model (left) and its RD curve (right).
$R=1.89,D=0.98$
$R=2.55,D=0.51$
Figure 5: Visualization of conditional sampling with different $(R,D)$ pairs
Figure 6: Comparison between the original distribution and the conditional sampling distribution at $R=5.96,D=0.02$. The blue scatter ones represent the original data, and the orange points represent the sampled data.
Appendix A Estimation of rate and distortion
[…]
^$\widehat{\mathcal{L}}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{1}{N}\sum_{j=1}^{N}e^{-\beta\rho(x_{i},y_{j})}\right).$ (5)
Further, the distortion is estimated by
$D$ $=\int_{\mathcal{X\times Y}}p(x)p(y|x)\rho(x,y)\mathrm{d}x\mathrm{d}y$
$=\int_{\mathcal{X}}p(x)\frac{\int_{\mathcal{Y}}q(y)e^{-\beta\rho(x,y)}\rho(x,y)\mathrm{d}y}{\int_{\mathcal{Y}}q(y)e^{-\beta\rho(x,y)}\mathrm{d}y}\mathrm{d}x$
$\approx\frac{1}{N}\sum_{i=1}^{N}\frac{\sum_{j=1}^{N}e^{-\beta\rho(x_{i},y_{j})}\rho(x_{i},y_{j})}{\sum_{j=1}^{N}e^{-\beta\rho(x_{i},y_{j})}}.$
Then, the rate is obtained by
$R=\widehat{\mathcal{L}}(\theta)-\beta D.$
Next, we analyze the asymptotic relationship between the estimated objective function (5) and the original objective function (2), which is precisely described by the following theorem.
Theorem 1.
Suppose that there exists $M$, such that for $\forall x\in\mathcal{X}$, it satisfies that $\int_{\mathcal{Y}}e^{-\beta\rho(x,y)}\mathrm{d}y<M$. Let $F^{*}$ be the optimal value of (2), and then there exists a compact parameter domain $\Theta$ such that the optimal value of ^$\widehat{\mathcal{L}}(\theta)$, defined as
$\widehat{F}_{N}=\inf_{\theta\in\Theta}\widehat{\mathcal{L}}(\theta),$
converges to $F^{*}$ with probability one. Specifically, for $\forall\epsilon>0$, there exist a positive integer $N_{0}$, such that $\forall N>N_{0}$,
$|\widehat{F}_{N}-F^{*}|<\varepsilon,\quad a.e.$ (6)
where “a.e.” stands for almost everywhere.
Proof.
Fix $\varepsilon>0$. Let $q^{*}(y)$ be the optimal solution that minimizes (2), and it takes the form of the Boltzmann distribution:
$q^{*}(y)=\frac{e^{-E^{*}(y)}}{Z^{*}}.$
where $Z^{*}=\int_{\mathcal{Y}}e^{-E^{*}(y)}\mathrm{d}y$ and $E^{*}(y)>0$.
Fix $\xi>0$. By the universal approximation theorem [24, 25], there exist $\Theta$ and $\hat{\theta}\in\Theta$, such that $\forall y\in\mathcal{Y}$,
$|E^{*}(y)-E_{\hat{\theta}}(y)|<\xi.$
Then it holds that
$|e^{-E^{*}(y)}-e^{-E_{\hat{\theta}}(y)}|$ $=e^{-E^{*}(y)}|1-e^{E^{*}(y)-E_{\hat{\theta}}(y)}|$
$<e^{-E^{*}(y)}(e^{\xi}-1).$
From this, one obtains
$|q^{*}(y)-q^{\hat{\theta}}(y)|=|\frac{e^{-E^{*}(y)}}{Z^{*}}-\frac{e^{-E_{\hat{\theta}}(y)}}{Z_{\hat{\theta}}}|$
$\leq$ $\frac{e^{-E^{*}(y)}|Z_{\hat{\theta}}-Z^{*}|+Z^{*}|e^{-E^{*}(y)}-e^{-E_{\hat{\theta}}(y)}|}{Z^{*}Z_{\hat{\theta}}}$
$<$ $\frac{2(e^{\xi}-1)}{(2-e^{\xi})Z^{*}}.$
This further implies that
$\left|\int_{\mathcal{Y}}q^{*}(y)e^{-\beta\rho(x,y)}\mathrm{d}y-\int_{\mathcal{Y}}q_{\hat{\theta}}(y)e^{-\beta\rho(x,y)}\mathrm{d}y\right|$
$<$ $M\frac{2(e^{\xi}-1)}{(2-e^{\xi})Z^{*}}.$
Since the right-hand side of the above inequality tends to zero as $\xi\rightarrow 0$ and $\log x$ is continuous, there exists a positive number $\xi$, such that the following relation holds:
$\left|\log\int_{\mathcal{Y}}q^{*}(y)e^{-\beta\rho(x,y)}\mathrm{d}y-\log\int_{\mathcal{Y}}q_{\hat{\theta}}(y)e^{-\beta\rho(x,y)}\mathrm{d}y\right|<\frac{\varepsilon}{2}.$
This means that by choosing a suitable $\hat{\theta}$, one has $|F^{*}-\mathcal{L}(\hat{\theta})|<\frac{\varepsilon}{2}$. Next, define
$F_{\Theta}=\inf_{\theta\in\Theta}\mathcal{L}(\hat{\theta}).$
From this definition, one obtains $F^{*}\leq F_{\Theta}\leq\mathcal{L}(\hat{\theta})$, and then
$|F^{*}-F_{\Theta}|<\frac{\varepsilon}{2}.$ (7)
Further, one has
$|\widehat{F}_{N}-F_{\Theta}|=|\inf_{\theta\in\Theta}\widehat{\mathcal{L}}(\theta)-\inf_{\theta\in\Theta}\mathcal{L}(\theta)|\leq\sup_{\theta\in\Theta}|\widehat{\mathcal{L}}(\theta)-\mathcal{L}(\theta)|.$
It follows from the law of large numbers [26] that for any given $\epsilon>0$, there exists $N_{0}$, such that $\forall N>N_{0}$ and with probability one, |^$|\widehat{\mathcal{L}}(\theta)-\mathcal{L}(\theta)|\leq\frac{\varepsilon}{2}$. Therefore
$|\widehat{F}_{N}-F_{\Theta}|<\frac{\varepsilon}{2}.$ (8)
Combining (7) and (8), one finally has
$|F^{*}-\widehat{F}_{N}|<\varepsilon,\,\forall\,N>N_{0}.$
This concludes the proof of the theorem. ∎